OpenStudy (elise_a18):

GEOMETRIC SEQUENCES The original value of a car is $18,000, and it depreciates (loses value) by 15% each year. What is the value of the car after three years?

OpenStudy (eliesaab):

After one year it loses .15 *18000 So how much is it worth after one year

OpenStudy (elise_a18):

12000? I have no idea

OpenStudy (eliesaab):

18000 - .15 18000=18000 -2700

OpenStudy (elise_a18):

15300

OpenStudy (eliesaab):

18000-2700=15300

OpenStudy (eliesaab):

Yes

OpenStudy (elise_a18):

that doesn't help too much for figuring out 3 years

OpenStudy (eliesaab):

Do another year but strat with 15300

OpenStudy (elise_a18):

strat?

OpenStudy (eliesaab):

15300 - .15 * 15300

OpenStudy (eliesaab):

How much is that?

OpenStudy (elise_a18):

is there no formula to use where I can plug in numbers?

OpenStudy (eliesaab):

Yes, I will get you to the formula when you follow my steps

OpenStudy (elise_a18):

isn't that just making the process longer? You could give me the formula to use and I would know what to do from there :)

OpenStudy (eliesaab):

\[18000*(1-.15)^3\]

OpenStudy (elise_a18):

17939.25

OpenStudy (eliesaab):

No

OpenStudy (elise_a18):

what did i do incorrect? ono

OpenStudy (eliesaab):

\[ 18000 *(0.85)^3=11054.2 \]

OpenStudy (elise_a18):

ahh i see what i did... but why is the exponent on the outside?

OpenStudy (eliesaab):

The exponent is only for .85

OpenStudy (eliesaab):

18000 is not raised to the power 3

OpenStudy (elise_a18):

ahhh I HATE MAKING STUPID LITTLE MISTAKES

OpenStudy (elise_a18):

srry about caps... X)

OpenStudy (eliesaab):

In general if the original price is P and the depreciation is r %, then the value after in years is \[ P*(1-r)^n \]

OpenStudy (eliesaab):

Do not worry

OpenStudy (elise_a18):

thanks! i wouldn't have caught my mistake

OpenStudy (eliesaab):

NP. Love to help you/

OpenStudy (elise_a18):

i have one more if you would help..

OpenStudy (eliesaab):

Post is as new problem